The ratio of the interior angle to the exterior angle of a regular polygon is 3:2. Find the number of side In the polygon.


1
Expert's answer
2021-06-14T15:25:30-0400

Let "n=" the nu ber of sides of a regular polygon.

The measure of each interior angle is "\dfrac{(n-2)180\degree}{n}" and the measure of each exterior angle is "\dfrac{360\degree}{n}."

Given that the ratio of the interior angle to the exterior angle of a regular polygon is 3:2


"\dfrac{\dfrac{(n-2)180\degree}{n}}{\dfrac{360\degree}{n}.}=\dfrac{3}{2}"

"2(n-2)=3(2)"

"n=5"